1. Landon Lowman, star quarterback of the university football team, is thinking about forgoing his last two years of eligibility and making himself available for the professional football draft. Scouts estimate that Landon could receive a signing bonus of $1.5 million today along with a five-year contract for $3 million per year (payable at the end of the year). They further estimate that he could negotiate a contract for $5.5 million per year for the remaining seven years of his career.

The scout believe, however, that Landon will be a much higher draft pick if he improves by playing out his eligibility. If he stays at the university, he is expected to receive a $2.5 million signing bonus in two years along with a 5-year contract for $5 million per year. After that, the scouts expect Landon to obtain a five-year contract for $6 million per year to take him into retirement.

-

Assume that Landon can earn a 8% return over this time. Should Landon stay or go?

2. Suppose you make monthly mortgage payments of $1,345 and have 11 years left on the mortgage (next payment due next month). Assuming a 4.5% stated annual interest rate for the mortgage, how much would you need today to pay off the mortgage?
Answer #1

(1) If Landon Stays: Signing Bonus $ 2.5 million at the end of Year 2, Annual Payouts = $ 5 million for 5 years (at the end of Year 3, Year 4, Year 5, Year 6 and Year 7), Final contract payouts = $ 6 million (at the end of Year 8 to Year 12)

Interest Rate = 8%

Present Value of Cash Flows = 2.5 / (1.08)^(2) + 5 x (1/0.08) x [1-{1/(1.08)^(5)}] x [1/(1.08)^(2)] + 6 x (1/0.08) x [1-{1/(1.08)^(5)}] x [1/(1.08)^(7)] = $ 33.237 million ~ $ 33.24 million

If Landon Leaves: Upfront Payment of $ 1.5 million, Annual Payouts = $ 3 million for 5 years and Final Payout = $ 5.5 million for the remaining 7 Years

Interest Rate = 8 %

PV of Cash Flows = 1.5 + 3 x (1/0.08) x [1-{1/(1.08)^(5)}] + 5.5 x (1/0.08) x [1-{1/(1.08)^(7)}] x [1/(1.08)^(5)] = $ 32.97 million

As the PV of Cash Flows is higher in case Landon Stays, Landon should stay instead of leaving.

NOTE: Please raise a separate query for the solution to the second unrelated question as one query is restricted to the solution of only one complete question with up to four sub-parts.